OpenStudy (anonymous):

Determine whether the point (8, 6) is a solution of the system 11x – 8y = 40 -6x + 1y = -42 by substituting the point into each equation. Determine whether the point (8, 6) is a solution of the system 11x – 8y = 40 -6x + 1y = -42 by substituting the point into each equation. @Mathematics

OpenStudy (anonymous):

88-48 = 40 this is correct -48+6 = 42 correct yes it is

jimthompson5910 (jim_thompson5910):

11x – 8y = 40 11(8) – 8(6) = 40 88-48 = 40 40 = 40 ... works ------------------- -6x + 1y = -42 -6(8) + 1(6) = -42 -48 + 6 = -42 -42 = -42 ... works Since both equations are satisfied, the ordered pair is indeed a solution.

OpenStudy (anonymous):

okay i see where i made my mistakes thanks guys

OpenStudy (anonymous):

i was not multiplying